# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ 3 Identification and Bounds
$L(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{X\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{=}(X)]$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{<}(X)]$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{\not\leq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{\text{min}}$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X,D\in S_{*}(X)]$
$\quad+\mathbf{1}\{X\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{\text{min}}]$
$U(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{X\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{=}(X)]$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{>}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{>}(X)]$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{\not\geq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{\text{max}}$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{*}(X)]$
$\quad+\mathbf{1}\{X\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{\text{max}}]$
These bounds are still valid if we replace $\mathcal{S}_{<}(x)$ with $\tilde{\mathcal{S}}_{<}(x)$ and $\mathcal{S}_{>}(x)$ with $\tilde{\mathcal{S}}_{>}(x)$, which we denote $\tilde{L}(\pi_{e}),\tilde{U}(\pi_{e})$. We give intuition for the construction of the lower bound. First, under model output $a_{0}$, we use data from the control arm and any trialed model outputting $a_{0}$ to infer patient outcomes. Next, we break up the target estimand into its components over three subpopulations: two subpopulations where there is at least one trialed model in the RCT that matches in output with $\pi_{e}$ and that (i) has the same performance as $\pi_{e}$ and (ii) has worse performance than $\pi_{e}$, and a third subpopulation that fits into neither of the above two categories. In the first subpopulation, we point identify the policy value from data trialing models that match in output and have the same performance. In the second subpopulation, we lower bound the policy value with data trialing models that match in output and have worse performance. Finally, in the third subpopulation, we have no information on how $\pi_{e}$ might behave, so we lower bound the policy value by $Y_{\text{min}}$. Notably, these bounds cannot be improved without further assumptions, as we show in the following.
@@
[Tightness of bounds in Assumption 3.3]theoremNonIdBounds For any observational distribution $P(X,Y,A,M,\Pi,D)$ consistent with the assumptions of Assumption 3.3, there exist two structural causal models $\mathcal{M}_{L},\mathcal{M}_{U}$ such that both are consistent with Assumptions 2.1, 3.1, 3.2 and 3.3, both give rise to that same observational distribution $P$, and where the policy value of any policy $\pi_{e}$ under $\mathcal{M}_{L},\mathcal{M}_{U}$ is given by $\tilde{L}(\pi_{e}),\tilde{U}(\pi_{e})$ from Assumption 3.3, respectively (the bounds constructed using $\tilde{\mathcal{S}}_{<},\tilde{\mathcal{S}}_{>}$). Hence, these bounds cannot be improved without further assumptions. Note that Assumption 3.3 implies that the tightest possible bounds require the use of e.g., $\tilde{\mathcal{S}}_{<}$ in place of $\mathcal{S}_{<}$. This requirement arises because we may get tighter lower-bounds (and similarly, tighter upper-bounds) by using only outcomes under the “next-worst/best” performing model, rather than averaging over outcomes under all worse/better-performing models. Nonetheless, it may be useful to use $\mathcal{S}$ instead of $\tilde{\mathcal{S}}$ in some scenarios due to sample-size concerns, especially if outcomes $Y$ do not vary significantly with $M$.
When is exact identification possible? To better understand conditions for agreement of upper and lower bounds, we directly consider the width of these bounds. {restatable}[Bound Decomposition]propositionBoundDecomposition The gap between the bounds in Assumption 3.3 can be written as
$U(\pi_{e})-L(\pi_{e})$ $=\mathbb{E}[\delta(X,Y,D)]$
@@
$\delta(X,Y,D)=(Y_{\text{max}}-Y_{\text{min}})\mathbf{1}\{X\in\mathcal{X}_{*}^{c}\}$ (4)
$+(\mathbb{E}[Y\mid X,D\in\mathcal{S}_{>}(X)]-\mathbb{E}[Y\mid X,D\in\mathcal{S}_{<}(X)])$
$\quad\cdot\mathbf{1}\{\pi_{e}\neq a_{0},X\in\mathcal{X}_{<}\cap\mathcal{X}_{>}\}$ (5)
$+(Y_{\text{max}}-\mathbb{E}[Y\mid X,D\in\mathcal{S}_{<}(X)])$
$\quad\cdot\mathbf{1}\{\pi_{e}\neq a_{0},X\in\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}\}$ (6)
$+(\mathbb{E}[Y\mid X,D\in\mathcal{S}_{>}(X)]-Y_{\text{min}})$
$\quad\cdot\mathbf{1}\{\pi_{e}\neq a_{0},X\in\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}\}$ (7)
Moreover, $\delta(X,Y,D)\geq 0$ almost surely under the assumptions of Assumption 3.3. For $\tilde{L},\tilde{U}$, this gap holds with $\mathcal{S}_{<}(x),\mathcal{S}_{>}(x)$ replaced with $\tilde{\mathcal{S}}_{<}(x),\tilde{\mathcal{S}}_{>}(x)$.
@@
Note that the only way to achieve point identification (a gap of zero in Assumption 3.3) is if each component (Eqs. 4, 5, 6 and 7) is equal to zero. Eq. 4 captures uncertainty that arises in the subset of the population $\mathcal{X}_{*}^{c}$ where no trialed model agrees with the output of the model $\pi_{e}$. For this term to be zero, there must exist some trialed model that agrees with the action taken by $\pi_{e}$ for every value of $\mathcal{X}$. Second, Eq. 5 reflects differences in the bounds for the subpopulation where the evaluation policy $\pi_{e}$ takes a non-neutral action ($\pi_{e}(X)\neq a_{0}$), and the trialed models that agree with $\pi_{e}$ include both better-performing ($\mathcal{X}_{>}$) and worse-performing ($\mathcal{X}_{<}$) models [^6] Recall that the definition of $\mathcal{X}_{<}$ and $\mathcal{X}_{>}$ explicitly excludes any values of $X$ where there exists a model/policy that agrees with $\pi_{e}$ and has exactly the same performance.. Here, the outcomes under the better/worse-performing models give an upper/lower bound on the outcomes under $\pi_{e}$. For this term to be zero, either this subpopulation $(\mathcal{X}_{<}\cap\mathcal{X}_{>})$ must be empty, or the outcomes under the better and worse-performing models must coincide [^7] Equivalence of conditional outcomes could occur if differences in performance do not impact outcomes for the range of performances tested. Note that Assumption 3.1 allows for this occurrence, as it does not assume a strict inequality.. Equations 6 and 7 capture subpopulations where the only models that agree with $\pi_{e}$ are either better-performing (giving an upper bound, but no meaningful lower bound on performance) or worse-performing, and these sets must be empty for these terms to be zero.
@@
How can we estimate these bounds from data? Assumption 3.3 below implies a simple estimator that can be used to estimate bounds (and provide asymptotically valid confidence intervals) on the policy value for a new policy $\pi_{e}$ without the need for training auxiliary models.
@@
propositionEmpiricalEstimation The bounds in Assumption 3.3 can be written $L(\pi_{e})=\mathbb{E}[\psi_{L}(X,D,Y)]$ and $U(\pi_{e})=\mathbb{E}[\psi_{U}(X,D,Y)]$, where $\psi_{L}$ and $\psi_{U}$ are defined as follows
$\psi_{L}(X,D,Y)$$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{=}(X)\}}{p(D\in\mathcal{S}_{=}(X))},&\ \text{if }X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{<}(X)\}}{p(D\in\mathcal{S}_{<}(X))},&\ \text{if }X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\\ Y_{\text{min}},&\ \text{if }X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{*}(X)\}}{p(D\in\mathcal{S}_{*}(X))},&\ \text{if }X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\\ Y_{\text{min}},&\ \text{if }X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\end{cases}$
$\psi_{U}(X,D,Y)$$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{=}(X)\}}{p(D\in\mathcal{S}_{=}(X))},&\ \text{if }X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{>}(X)\}}{p(D\in\mathcal{S}_{>}(X))},&\ \text{if }X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\\ Y_{\text{max}},&\ \text{if }X\in\mathcal{X}_{\not\geq},\pi_{e}(X)\neq a_{0}\\ Y\cdot\frac{\mathbf{1}\{D\in\mathcal{S}_{*}(X)\}}{p(D\in\mathcal{S}_{*}(X))},&\ \text{if }X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\\ Y_{\text{max}},&\ \text{if }X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\\ \end{cases}$
@@
$\bigg{[}\hat{L}(\pi_{e})-\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\cdot\frac{\hat{\sigma}(\psi_{L})}{\sqrt{n}},\hat{U}(\pi_{e})+\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\cdot\frac{\hat{\sigma}(\psi_{U})}{\sqrt{n}}\bigg{]}$
@@
is an asymptotically valid ($1-\alpha$)-confidence interval, where $\hat{\sigma}(\psi)$ is the empirical standard deviation of $\psi$ and $\Phi^{-1}$ is the inverse of the standard normal CDF. For $\tilde{L},\tilde{U}$, all the above holds after replacing $\mathcal{S}_{<}(x),\mathcal{S}_{>}(x)$ with $\tilde{\mathcal{S}}_{<}(x),\tilde{\mathcal{S}}_{>}(x)$.
@@
Assumption 3.3 gives a straightforward way of estimating bounds from data as empirical means over the RCT dataset. To give intuition, terms like $p(D\in\mathcal{S}_{>}(X))$ are known by design (though they depend on $X$), since $p(D)$ is assumed to be known and the trialed policies $\Pi$ are known, and so asymptotic normality is straightforward to demonstrate. [^8] The statistical efficiency of these bounds could potentially be improved by using a doubly-robust-style estimator that incorporates an estimate of terms such as $\mu_{=}(X)\coloneqq E[Y\mid X=x,D\in\mathcal{S}_{<}(X)]$, but we present this simpler estimator for ease of exposition. In Section 5, we use Assumption 3.3 to estimate the bounds for the effect of deploying a new model.
4 Implications for pre-trial design and post-trial analysis
Aside: Why require the performance assumption? We pause to reflect on the importance of the assumption (implicit in Assumption 2.1) that implies that our choice of model $\Pi$ impacts outcomes, not only through the outputs $A$, but also through model performance $M$. Broadly speaking, this assumption is not only intuitive from a real-world perspective, but it also has the welcome side-effect of ruling out nonsensical conclusions about trial design. For instance, there are trivial ways to satisfy the condition that, for every value of $\mathcal{X}$, there exists some model in the trial that matches the output of $\pi_{e}$. In the context of Example 1, for instance, one could trial an alerting system that simply always raises alerts for every patient, alongside a control arm that never raises alerts. The assumption that model accuracy $M$ impacts outcomes gives a formal rationale for why this type of trial design is nonsensical: The observed impact of this “always alert” policy would likely be minimal (or even harmful), compared to never raising alerts, due to the negative impact of extremely poor accuracy or false positive rates.

### v2

@@ 3 Identification and Bounds
$L(\pi_{e})=\mathbb{E}\big{[}$ ${\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\mathbf{1}\{\pi_{e}\neq a_{0}\}}\big{(}$
${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)]$ (4)
$+$ ${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}}Y_{min}\big{)}$ (5)
$+$ ${\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\mathbf{1}\{\pi_{e}=a_{0}\}}\big{(}$
${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(X)]$ (6)
$+$ ${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}}Y_{min}\big{)}\big{]}$ (7)
$U(\pi_{e})=\mathbb{E}\big{[}$ ${\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\mathbf{1}\{\pi_{e}\neq a_{0}\}}\big{(}$
${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)]$
$+$ ${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}}Y_{max}\big{)}$
$+$ ${\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\mathbf{1}\{\pi_{e}=a_{0}\}}\big{(}$
${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(X)]$
$+$ ${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}}Y_{max}\big{)}]$
These bounds are still valid if we replace $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$ with $\mathbf{\Pi}^{e}_{\leq}(X)$ and $\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)$ with $\mathbf{\Pi}^{e}_{\geq}(X)$. We give intuition for the construction of the lower bound. First, note that each value of $x\in\mathcal{X}$ makes a contribution to the construction of the lower bound based on which green and which purple indicator it activates. In Eq. 4, we consider model outputs that are not the neutral action $a_{0}$ and values of $X$ at which $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$ is non-empty. That is, there is at least one trialed model agreeing in output with $\pi_{e}$ that also has less than or equal performance. Here, we use outcome data from trial arms with the next-worst performance to infer outcomes. Next, Eq. 5 represents values of $X$ where $\pi_{e}$ does not output the neutral action and there are no agreeing trialed models with less than or equal performance. In this case, we use $Y_{min}$ to lower bound outcomes as we have no trial data on outcomes under such model output. In Eq. 6, the new model outputs the neutral action $a_{0}$, and there is at least one trial model agreeing in output. We can then use data from any trial models that agreed in output to infer outcomes as model performance does not influence outcomes under the neutral action. Finally, Eq. 7 represents no trial models agreeing in an output of the neutral action; here, we lower bound by $Y_{min}$. The intuition for the upper bound follows similarly.
The lower and upper bounds in Remark 3.1 can be constructed by iterating over all possible values of $X$, determining whether the model output is a neutral action, checking whether there are agreeing trial models with appropriate levels of performance, and taking the weighted average of the appropriate bounds given the observations above over $X$. In Appendix B, we give an algorithm for constructing the bounds proposed in Remark 3.1 in this manner.
@@
[Tightness of bounds in Remark 3.1]theoremNonIdBounds For any observational distribution $P(X,Y,A,M,\Pi,D)$ consistent with the assumptions of Remark 3.1, there exist two structural causal models $\mathcal{M}_{L},\mathcal{M}_{U}$ such that both are consistent with Assumptions 2.1, 3.1, 3.2 and 3.3, both give rise to that same observational distribution $P$, and where the policy value of any policy $\pi_{e}$ under $\mathcal{M}_{L},\mathcal{M}_{U}$ is given by $L(\pi_{e}),U(\pi_{e})$ from Remark 3.1, respectively. Hence, these bounds cannot be improved without further assumptions. Note that Remark 3.1 implies that the tightest possible bounds require the use of $\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)$ and $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)$ in place of $\mathbf{\Pi}^{e}_{\leq}(x)$ and $\mathbf{\Pi}^{e}_{\geq}(x)$, respectively. This requirement arises because we may get tighter lower-bounds (and similarly, tighter upper-bounds) by using only outcomes under the “next-worst/best” performing model, rather than averaging over outcomes under all worse/better-performing models. Nonetheless, it may be useful to use $\mathbf{\Pi}^{e}_{\leq}(x)$ instead of $\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)$ in some scenarios due to sample-size concerns, especially if outcomes $Y$ do not vary significantly with $M$.
When is exact identification possible? To better understand conditions for agreement of upper and lower bounds, we directly consider the width of these bounds. {restatable}[Bound Decomposition]propositionBoundDecomposition The gap between the bounds in Remark 3.1 can be written as
$U(\pi_{e})-L(\pi_{e})$ $=\mathbb{E}[\delta(X,Y,\Pi)]$
@@
$\delta(X,Y,\Pi)=$
${\color[rgb]{0,0.46484375,0.734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.46484375,0.734375}\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}}(Y_{\text{max}}-Y_{\text{min}})$ (8)
$+$ ${\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\mathbf{1}\{\pi_{e}\neq a_{0}\}}\big{[}$
${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}}\cdot$ (9)
$(\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)]-\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)])$
$+$ ${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}}\cdot$ (10)
$(Y_{\text{max}}-\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)])$
$+$ ${\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}}\cdot$ (11)
$(\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)]-Y_{\text{min}})\big{]}$
Moreover, $\delta(X,Y,\Pi)\geq 0$ almost surely under the assumptions of Remark 3.1.
@@
Note that the only way to achieve point identification (a gap of zero in Remark 3.1) is if each component (Eqs. 8, 9, 10 and 11) is equal to zero. Here, a value of $x\in\mathcal{X}$ only contributes to the bound decomposition if $x$ satisfies a blue indicator or both a green and purple indicator. Eq. 8 captures uncertainty that arises in the subset of the population, indexed by $x\in\mathcal{X}$, for which $\{\mathbf{\Pi}^{e}(x)=\varnothing\}$ where no trialed model agrees with the output of the model $\pi_{e}$ regardless of whether the new model outputs $a_{0}$. For this term to be zero, there must exist some trialed model that agrees with the action taken by $\pi_{e}$ for every value of $x\in\mathcal{X}$. Second, Eq. 9 reflects differences in the bounds for the subpopulation where the evaluation policy $\pi_{e}$ takes a non-neutral action ($\pi_{e}(X)\neq a_{0}$), and the trialed models that agree with $\pi_{e}$ have at least one of equal performance [^7] When there exists an agreeing trial model with equal performance, both $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$ and $\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}$ are true according to their definitions. or include both better-performing and worse-performing models. Here, the outcomes under the better/worse-performing models give an upper/lower bound on the outcomes under $\pi_{e}$. For this term to be zero, either there exists an agreeing trial model with equal performance, this subpopulation $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)\neq\varnothing\}$ is empty, or the outcomes under the better and worse-performing models coincide [^8] Equivalence of conditional outcomes could occur if differences in performance do not impact outcomes for the range of performances tested. Note that Assumption 3.1 allows for this occurrence, as it does not assume a strict inequality.. Equations 10 and 11 capture subpopulations where the only models that agree with $\pi_{e}$ are either worse-performing (giving a lower bound, but no meaningful upper bound) or better-performing (giving an upper bound, but no meaningful lower bound) when the new model does not output $a_{0}$, and these sets must be empty for these terms to be zero.
@@
How can we estimate these bounds from data? Remark 3.1 below implies a simple estimator that can be used to estimate bounds (and provide asymptotically valid confidence intervals) on the policy value for a new policy $\pi_{e}$ without the need for training auxiliary models.
@@
propositionEmpiricalEstimation The bounds in Remark 3.1 can be written $L(\pi_{e})=\mathbb{E}[\psi_{L}(Y,X,\Pi)]$ and $U(\pi_{e})=\mathbb{E}[\psi_{U}(Y,X,\Pi)]$, where $\psi_{L}$ and $\psi_{U}$ are defined as follows
$\psi_{L}(Y,X,\Pi)$
$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\}}{P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X))},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)\neq a_{0}}\\ Y_{min},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)\neq a_{0}}\\ Y\cdot\frac{\mathbf{1}\{\Pi\in\mathbf{\Pi}^{e}(X)\}}{P(\Pi\in\mathbf{\Pi}^{e}(X))},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{\Pi}^{e}(X)\neq\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)=a_{0}}\\ Y_{min},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{\Pi}^{e}(X)=\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)=a_{0}}\\ \end{cases}$
$\psi_{U}(Y,X,\Pi)$
$\coloneqq\begin{cases}Y\cdot\frac{\mathbf{1}\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\}}{P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X))},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)\neq a_{0}}\\ Y_{max},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)\neq a_{0}}\\ Y\cdot\frac{\mathbf{1}\{\Pi\in\mathbf{\Pi}^{e}(X)\}}{P(\Pi\in\mathbf{\Pi}^{e}(X))},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{\Pi}^{e}(X)\neq\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)=a_{0}}\\ Y_{max},&\ \text{if }{\color[rgb]{0.66796875,0.265625,0.6015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.66796875,0.265625,0.6015625}\mathbf{\Pi}^{e}(X)=\varnothing},{\color[rgb]{0.06640625,0.46484375,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.06640625,0.46484375,0.19921875}\pi_{e}(X)=a_{0}}\\ \end{cases}$
@@
$\hat{U}(\pi_{e})+\Phi^{-1}\left(1-\frac{\alpha}{2}\right)\cdot\frac{\hat{\sigma}(\psi_{U})}{\sqrt{n}}\bigg{]}$
@@
is an asymptotically valid ($1-\alpha$)-confidence interval, where $\hat{\sigma}(\psi)$ is the empirical standard deviation of $\psi$ and $\Phi^{-1}$ is the inverse of the standard normal CDF.
@@
Remark 3.1 gives a straightforward way of estimating bounds from data as empirical means over the RCT dataset. To give intuition, terms like $P(\Pi\in\mathbf{\Pi}^{e}_{\geq}(X))$ are known by design (though they depend on $X$), since $P(\Pi)$ is assumed to be known and the trialed policies $\varPi$ are known, and so asymptotic normality is straightforward to demonstrate [^9] The statistical efficiency of these bounds could potentially be improved by using a doubly-robust-style estimator that incorporates an estimate of terms, such as $\mu_{=}(X)\coloneqq E[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)]$, but we present this simpler estimator for ease of exposition and understanding.. In Section 5, we use Remark 3.1 to estimate the bounds for the effect of deploying a new model.
4 Recommendations For Pre-trial Design